OpenStudy (anonymous):

Help needed with finding the slope intercept of a line. I don't understand why/how. (0,-2) m=3 y+2=3(x-0)

OpenStudy (anonymous):

y-y1=m(x-x1)---- this is the formula used to find a line where y1=-2 and x1=0

OpenStudy (anonymous):

now can specify what you dont understand

OpenStudy (anonymous):

yes I get the whole plug in numbers

OpenStudy (anonymous):

solve for y now

OpenStudy (anonymous):

Slope intercept is a form of the line in which you need the slope and the y-intercept. y=mx+b m=slope b=y-intercept The y-intercept occurs ant the point where 'x'=0 You are therefore given the y-intercept through (0,-2), and they plainly told you 'm' Just plug these into y=mx+b

OpenStudy (anonymous):

y+2=3x y=3x-2 then what?

OpenStudy (anonymous):

That's it.

OpenStudy (anonymous):

yeh thats is it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

y=mx+b y=3x+b b is negative 2, so y=3x-2